0.5 g of a metallic oxide is converter to 1.165 g of the corresponding sulphate. If the atomic mass of the element is 88g/mol, the metallic oxide.is ??

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
the metallic oxide.is
we have
MOx --------------->M(SO4)x
applying principle of atomic conservation,we get
2*.5/(88*2+16x) =2*1.165/(88*2 +96x)
.5/(176+16x) =1.165/(176+96x
solving for x,we get x=4
Therefore the metal oxide is MO4